a stone is dropped from the top of a 45m high building how fast will it be moving When it reaches the ground and what will its velocity be?


1
Expert's answer
2021-03-29T08:58:25-0400

Given


H = 45m

and we take g(acceleration due to gravity )=9.8m/s2

And

V(velocity at ground)


By using energy conservation

or

using Newton's equation of motion

{v2- u2 =2as }

here u (initial velocity) = 0

v=v

a = g

s = h

By solving , we get


V = "\sqrt{2gh}"


"\implies" V= "\sqrt {2\times 9.8\times 45}"


"\implies" v="\sqrt{882}"


"\implies" "\boxed{v=29.698m/s}" Towards the ground